One-fourth of a two digit number is three more than one-seventh of the number. What is the sum of the digits of that number ?

Answer by checkley79(3341)

X/4=X/7+3
X/4-X/7=3
(7X-4X)/4*7)=3
3X/28=3 CROSS MULTIPLY.
3X=28*3
3X=84
X=84/3
X=28 IS THE NUMBER.
PROOF:
28/4=28/7+3
7=4+3
7=7
2+8=10 ANS. FOR THE SUM OF THE DIGITS.